Suppose two coins are tossed. Let B be the random variable representing the number of heads that occur find the probability of getting single head.



1
Expert's answer
2022-03-28T12:53:12-0400

P(B=1)=P(H,T)+P(T,H)=(1/2)(1/2)+(1/2)(1/2)=2/4=0.5